OpenStudy (anonymous):

I don't get it... using A = P(1 + R)n. 6. Suppose that you invest $850 in an account that earns interest at an APR of 5.9%, compounded quarterly. Determine the accumulated balance after 22 years

OpenStudy (anonymous):

22yrs = 88quarters

OpenStudy (anonymous):

r= 5.9% =0.059/year so rate per quarter = 0.059 / 4

OpenStudy (anonymous):

then sub in \[A = 850 ( 1+ \frac{0.059}{4} ) ^{88} \]

OpenStudy (anonymous):

then use a calculator to evaluate that, be careful with brackets you can evaluate it

OpenStudy (anonymous):

why would it be a power of ^88?

OpenStudy (anonymous):

power is no of full years or numbers of quarters 1 year has 4 quarters so 22 years hv 22x4 = 88 quarters so power becomes 88

OpenStudy (anonymous):

Ok.. got it...

OpenStudy (anonymous):

now... would the same formula be used for a saving plan? Such as... Your saving account pays an APR of 4.6%, compounded annually. If you deposit $1850 at the end of each year for 8 years, what will be the accumulated balance in the account?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

first depositi becomes 1850(1.046)^8 because its in the bank for 8 year second deposit becomes 1850(1.046)^7 etc... then add them up

OpenStudy (anonymous):

hmm.. I got $4,381.85.... that doesn't seem right...